# Density-Matrix Mean-Field Theory

### v1

@@ Abstract
Mean-field theories (MFTs) have proven to be efficient tools for exploring various phases of matter, complementing alternative methods that are more precise but also more computationally demanding. Conventional mean-field theories (MFTs) often fall short in capturing quantum fluctuations, which restricts their applicability to systems characterized by strong quantum fluctuations. In this article, we propose a novel mean-field theory, density-matrix mean-field theory (DMMFT). DMMFT constructs effective Hamiltonians, incorporating quantum environments shaped by entanglements quantified by the reduced density matrices. Therefore, it offers a systematic and unbiased approach to account for effects of fluctuations and entanglements in quantum ordered phases. As demonstrative examples, we show that DMMFT can not only quantitatively evaluate the renormalization of order parameters induced by quantum fluctuations but can even detect the topological order of quantum phases. Additionally, we discuss the extensions of DMMFT for systems at finite temperatures and those with disorders. Our work provides a novel and efficient approach to explore phases exhibiting unconventional quantum orders, which can be particularly beneficial for investigating frustrated spin systems in high spatial dimensions.
@@ 1 Introduction
Exact approaches frequently encounter limitations posed by the exponential wall. In exact diagonalizations (ED), the dimension of the Hilbert space increases exponentially with the system size. In the density-matrix renormalization group (DMRG) [11], the area law provides relief from the exponential wall issue for gapped systems in one spatial dimension ($d=1$), it still faces exponential scaling challenges for gapless systems or in higher dimensions ($d\geq 2$)[12, 13, 14]. On the other hand, quantum Monte Carlo (QMC) methods are less constrained by the system size. Nevertheless, the notorious sign problem often plagues fermionic systems and frustrated magnetic systems [15, 16].
Approximation methods, serving as complements to exact approaches, prove to be useful and efficient tools for exploring various phases of the systems. The usual mean-field theories (MFTs) already provide insights into non-trivial effects arising resulting from the interactions, e.g., the formation of the local moments in metals [17], and the BCS theory for the superconductivity [18]. Conventional MFTs achieve the simplification by neglecting the fluctuations, Consequently, they tend to exhibits bias towards ordered states and overlook the nuanced effects stemming from the fluctuations.
@@
Despite of the successes of DMFT and its simplifications for fermionic systems, a gap remains in methods beyond semiclassical mean-field approximations for spin systems. Although employing the Holstein-Primakoff transformation [26] allows spins to be mapped to bosons, to which the DMFT may be adapted in principle, the transformation is nonlinear, and the semicalssical large-$S$ expansion becomes less controllable for $S=1/2$ in the quantum limit, posing challenges for semiclassical methods applied to quantum spins. Similar challenges may be encountered with alternative methods. For instance, in the Schwinger boson representation [27], one needs to set $\mathcal{N}=2$ for the quantum limit after a saddle-point mean-field approximation with a large $\mathcal{N}$
@@
In this article, we propose a new mean-field method beyond conventional MFTs, which we call density-matrix mean-field theory (DMMFT). DMMFT constructs effective Hamiltonians, incorporating quantum environments shaped by entanglements quantified by the reduced density matrices without presumed semiclassical orders. Therefore, it offers an unbiased approach to account for effects of fluctuations and entanglements in quantum ordered phases. In contrast to QMC and DMFT, DMMFT is generically applicable to systems of fermions, bosons as well as spins, regardless of frustrations. More importantly, by gauging the quantum fluctuations with the reduced DM, DMMFT can detect not only the symmetry-breaking phases in Landau’s paradigm but also the topological phases with the help of entanglement spectra. Our work provides a novel and efficient approach to explore phases exhibiting unconventional quantum orders, Particularly, it fills the gap left by the MFTs in studying quantum ordered phases in frustrated spin systems, where semiclassical methods become less controllable in the quantum limit, and QMC methods fail due to the sign problem.
@@
The rest of the article is organized as follows. In Sec. 2, we present the generic formulation of the DMMFT. The mean-field equations of DMMFT are derived in parallel to those of conventional MFTs. Moreover, we demonstrate that DMMFT becomes equivalent to conventional MFTs when the hyperparameter, which gauge the quantum fluctuations, is minimized. In Sec. 3, we apply the DMMFT to two demonstrative examples, 1) the Affleck-Khomoto-Lieb-Tasaki (AKLT) model [28, 29, 30], and 2) the antiferromagnetic Heisenberg model on triangular lattices (AFHTL) [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. In the AKLT model, DMFT identifies the topological ground states through their entanglement spectra. In the AFHTL, DMMFT reveals that quantum fluctuations not only renormalize the order parameters but also shift the phase boundaries. In Sec. 4, we compare the DMMFT with the DMRG and the DMFT. Additionally, we discuss the extensions of DMMFT for systems at finite temperatures and those with disorders. Finally, in Sec. 5, we draw conclusions and discuss potential avenues for future research
@@ 2 Density-Matrix Mean-Field Theory
Now, we formulate the DMMFT for the Hamiltonian in Eq.(1) Since no further assumptions are made specifying the local operators or the microscopic details of the coupling terms, DMMFT is generically applicable to fermions, bosons as well as spins, irrespective of the presence of frustration. In conventional MFTs, a local cluster is separated from the system and coupled to an effective environment, where the environment is assumed to be classical, and the correlated fluctuations between the cluster and the environment are neglected. DMMFT improves the conventional mean-field approximation by including the essential quantum fluctuations in the environment, where the reduced DM is used to gauge the quantum fluctuations and select a Hilbert subspaces approximating the effective environment.
In the following subsections, we will begin by reviewing the conventional mean-field approximation in Sec. 2.1. Then, we will develop the mean-field equations and self-consistency conditions of DMMFT in parallel to those in conventional MFTs in Sec. 2.2. Following this, in Sec. 2.3, we will conduct a comprehensive comparison between the DMMFT and the conventional MFTs. Within this comparison, we will identify a hyperparameter, $n_{c}$, that interpolates the DMMFT and the conventional MFTs. Particularly, when $n_{c}$ attains its minimal value of $1$, DMMFT becomes equivalent to the conventional MFTs.
@@ 2.1 Conventional Mean-Field Approximation
then, for each extended cluster, $|\phi_{\mathcal{C}^{\prime}_{c}}\rangle=\otimes_{c^{\prime}\in\mathcal{C}^{\prime}_{c}}|\phi_{c^{\prime}}\rangle$. Therefore, Eq.(4), Eq.(5), and Eq.(6) form a closed set of coupled mean-field equations for the conventional MFT.
The mean-field equations can often be further simplified when systems have additional symmetries. If the partition $\mathcal{C}=\{c\}$ respects some symmetry of the system, there exist symmetry transformations relating the clusters $T_{c^{\prime},c}:c\mapsto c^{\prime}$. A symmetric mean-field Ansatz state can be constructed simply as $|\Phi_{\text{MF}}\rangle=\otimes_{c}T_{c,c_{0}}|\phi_{c_{0}}\rangle$. Particularly, for a translationally invariant state, all $|\phi_{c}\rangle$ can be chosen equal to the same state $|\phi_{c_{0}}\rangle$ in the local Hilbert space $\mathscr{H}_{c_{0}}$, then the coupled mean-field equations for the clusters reduce to a single set of self-consistent mean-field equations for $|\phi_{c_{0}}\rangle$.
@@ 2.2 Density-Matrix Mean-Field Approximation
The entanglement is controlled by $\mathscr{H}_{c}$ than $\mathscr{H}_{\mathcal{C}\backslash c}$ provided $D_{\mathscr{H}_{c}}<D_{\mathscr{H}_{\mathcal{C}\backslash c}}$, where $D_{\mathscr{H}}$ denotes the dimension of a Hilbert space $\mathscr{H}$. According to the purification theorem [48], there exists a state $|\tilde{\Psi}\rangle$ in ~$\tilde{\mathscr{H}}=\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{\tilde{c}}$ such that $\text{Tr}_{\tilde{c}}|\tilde{\Psi}\rangle\langle\tilde{\Psi}|=\tilde{\rho}_{c}$ is equivalent to $\rho_{c}$, and the dimension of the Hilbert subspace ~$\tilde{\mathscr{H}}_{\tilde{c}}$ is bounded by $\tilde{D}^{E}_{c}=\lceil\exp({SE}_{c})\rceil$, where
@@
Instead of Eq.(4), DMMFT assumes that for systems without long-range entanglements, $-~\tilde{\mathscr{H}}$ can be found as a Hilbert subspace of $\mathscr{H}_{\bar{c}}$ over the extended cluster $\bar{c}$. Mathematically, there exists a homomorphism
@@
Parallel to the mean-field equations for the conventional MFTs, Eqs.(9) – (14) constitute a closed set of coupled mean-field equations for the DMMFT. The procedures for implementing the DMMFT algorithm are outlined in Appendix A. In the presence of symmetries, the mean-field equations can be further simplified. Specifically, for a translationally invariant target state $|\Phi\rangle$, the reduced DMs of the local clusters are all identical to the same reduced DM $\rho_{c_{0}}$ (as are the projectors). Consequently, the coupled mean-field equations for the clusters reduce to a single set of self-consistent mean-field equations for $\rho_{c_{0}}$.
@@ 2.3 Comparison of the Mean-Field Approximations
Comparing conventional MFTs with DMMFT, we find the cut-off parameter $n_{c}$ in Eq.(13) can be regarded as a hyperparameter interpolating the DMMFT and the conventional MFT. To demonstrate this point, we first observe that when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, DMMFT is equivalent to a conventional MFT. If the ground state is a product state [Eq.(6)], the local reduced density matrix of cluster $c$ is $\rho_{c}=|\phi_{c}\rangle\langle\phi_{c}|$. The expectation values of the local observables agree
@@
It is easy to observe that $\rho_{c}$ generically has non-vanishing entanglement entropy whenever $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}>1$. This is exactly the short-range entanglement captured by the DMMFT beyond the conventional MFTs. Not surprisingly, one expects that the DMMFT may detect short-range entangled topological phases with a properly chosen $n_{c}$. In Sec. 3.1, we demonstrate with the AKLT model that DMMFT can correctly detect the topological ground states in terms of the entanglement spectra without being biased to the symmetry-breaking states.
@@ 3 Applications
In this section, we apply the DMMFT to two demonstrative examples, the AKLT model and the AFHTL. These two models have been extensively studied in the literatures (cf. Ref. 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47 and the references therein). The AKLT model is an exactly solvable model with topological ground states, despiet the vanishing semiclassical (symmetry breaking) order parameters [29, 30]. On the other hand, the AFHTL represents a prototypical model for frustrated magnetism. With only nearest neighbor couplings and easy axis anisotropy, the AFHTL is known to order in a three-sublattice ordered state homologous to the semiclassical Néel state. However, the quantum fluctuations play non-trivial roles in stabilizing the ordered state [45, 46, 47] and significantly renormalize the order parameter [31]. In Sec. 3.1, we demonstrate that DMMFT can correctly detects the topological ground states based on the entanglement spectra. In Sec. 3.2, we compare the phases of the AFHTL in an external magnetic field determined by DMMFT and conventional MFT.

### v2

@@ Abstract
Mean-field theories have proven to be efficient tools for exploring various phases of matter, complementing alternative methods that are more precise but also more computationally demanding. Conventional mean-field theories often fall short in capturing quantum fluctuations, which restricts their applicability to systems characterized by strong quantum fluctuations. In this article, we propose a novel mean-field theory, density-matrix mean-field theory (DMMFT). DMMFT constructs effective Hamiltonians, incorporating quantum environments shaped by entanglements quantified by the reduced density matrices. Therefore, it offers a systematic and unbiased approach to account for effects of fluctuations and entanglements in quantum ordered phases. As demonstrative examples, we show that DMMFT can not only quantitatively evaluate the renormalization of order parameters induced by quantum fluctuations but can even detect the topological order of quantum phases. Additionally, we discuss the extensions of DMMFT for systems at finite temperatures and those with disorders. Our work provides a novel and efficient approach to explore phases exhibiting unconventional quantum orders, which can be particularly beneficial for investigating frustrated spin systems in high spatial dimensions.
@@ 1 Introduction
Exact approaches frequently encounter limitations posed by the exponential wall. In exact diagonalization (ED), the dimension of the Hilbert space increases exponentially with the system size. In density-matrix renormalization group (DMRG) [11], the area law provides relief from the exponential wall issue for gapped systems in one spatial dimension ($d=1$), it still faces exponential scaling challenges for gapless systems or in higher dimensions ($d\geq 2$)[12, 13, 14]. On the other hand, quantum Monte Carlo (QMC) methods are less constrained by the system size. Nevertheless, the notorious sign problem often plagues fermionic systems and frustrated magnetic systems [15, 16].
Approximation methods, serving as complements to exact approaches, prove to be useful and efficient tools for exploring various phases of a system. Conventional mean-field theories (MFTs) already provide insights into non-trivial effects arising from the interactions, e.g., the formation of the local moments in metals [17], and the BCS theory for the superconductivity [18]. Conventional MFTs achieve the simplification by neglecting the fluctuations, consequently, they tend to exhibit bias towards ordered states and overlook the nuanced effects stemming from the fluctuations.
@@
Despite of the successes of DMFT and its simplifications for fermionic systems, a gap remains in methods beyond semiclassical mean-field approximations for spin systems. Although employing the Holstein-Primakoff transformation [26] allows spins to be mapped to bosons, to which the DMFT may be adapted in principle, the transformation is nonlinear, and the semicalssical large-$S$ expansion becomes less controllable for $S=1/2$ in the quantum limit, posing challenges for semiclassical methods applied to quantum spins. Similar challenges may be encountered with alternative methods. For instance, in Schwinger boson representation [27], one needs to set $\mathcal{N}=2$ for the quantum limit after a saddle-point mean-field approximation with large $\mathcal{N}$
@@
In this article, we propose a new mean-field method beyond conventional MFTs, which we call density-matrix mean-field theory (DMMFT). DMMFT constructs effective Hamiltonians, incorporating quantum environments shaped by entanglements quantified by the reduced density matrices without presumed semiclassical orders. Therefore, it offers an unbiased approach to account for effects of fluctuations and entanglements in quantum ordered phases. In contrast to QMC and DMFT, DMMFT is generically applicable to systems of fermions, bosons as well as spins, regardless of frustrations. More importantly, by gauging the quantum fluctuations with the reduced DM, DMMFT can detect not only the symmetry-breaking phases in Landau’s paradigm but also the topological phases with the help of entanglement spectra. Our work provides a novel and efficient approach to explore phases exhibiting unconventional quantum orders. Particularly, it fills the gap left by the MFTs in studying quantum ordered phases in frustrated spin systems, where semiclassical methods become less controllable in the quantum limit, and QMC methods fail due to the sign problem.
Regarding DMMFT as a generalized cluster MFT, it is also worth noticing that there are alternative cluster based methods used for studying quantum spin systems, such as the cluster variation method (CVM) and the linked-cluster expansion (LCE). Following Kikuchi’s early study of the Ising model using [28], the CVM has also been extended for studying the quantum Heisenberg models [29, 30, 31]. CVM decomposes the entropic contribution to the free energy in terms of the cluster entropies that depends only on the reduced DM over the clusters. It becomes precise if the clusters up to the size of the entire system are considered [32]. CVM provides a systematic framework analyzing various cluster mean-field approximations as demonstrated in Ref. 29. Examined from the perspective of CVM, DMMFT includes the high-order cluster corrections containing essential information of quantum fluctuations beyond conventional MFTs. Similarly, LCE also decomposes the corrections to the physical observables into a series of terms depending on the clusters [33, 34, 35, 36, 37, 38]. An early development of the LCE treats the expansion of the free energy perturbatively in temperature, therefore it only performs well at high temperatures [33, 34, 35]. Improvement has been made by extending LCE for calculating cluster physical observables unperturbatively in temperature, also known as numerical LCE [36, 37, 38], which has been applied to some interesting frustrated quantum magnets [39, 40]. Complementing LCE below the critical temperature where long-range order develops, DMMFT does not suffer from convergence problems and can be more computationally efficient, as it does not require summation over a large number of clusters.
@@
The rest of the article is organized as follows. In Sec. 2, we present the generic formulation of the DMMFT. The mean-field equations of DMMFT are derived in parallel to those of conventional MFTs. Moreover, we demonstrate that DMMFT becomes equivalent to conventional MFTs when the hyperparameter that gauges the quantum fluctuations is minimized. In Sec. 3, we apply the DMMFT to two demonstrative examples, 1) the Affleck-Khomoto-Lieb-Tasaki (AKLT) model [41, 42, 43], and 2) the antiferromagnetic Heisenberg model on triangular lattices (AFHTL) [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. (Cf. Ref. 50 and references therein for a more comprehensive introduction to the frustrated Heisenberg models.) In the AKLT model, DMMFT is able to identify the topological ground states through their entanglement spectra. In the AFHTL, DMMFT reveals that quantum fluctuations not only renormalize the order parameters but also shift the phase boundaries. In Sec. 4, we compare DMMFT with DMRG and DMFT, as well as the cluster-based methods, CVM and LCE. Additionally, we discuss the extensions of DMMFT for systems at finite temperatures and those with disorders. In Sec. 5, we draw conclusions and discuss potential avenues for future research. To aid readers in understanding and applying DMMFT, appendices formulating the iterative algorithm for solving DMMFT equations and the pseudocodes implementing DMMFT for the AFHTL are presented in Append. A and B, respectively.
@@ 2 Density-Matrix Mean-Field Theory
Conventional MFTs isolate a local cluster from the system and coupled to an effective environment, where the environment is assumed to be classical, and the correlated fluctuations between the cluster and the environment are neglected. DMMFT improves the conventional mean-field approximation by including the essential quantum fluctuations in the environment, where the reduced DM is used to gauge the quantum fluctuations and select a Hilbert subspaces approximating the effective environment. In this section, we shall formulate DMMFT for the Hamiltonian in Eq.(1) without further assumptions specifying the local operators or the microscopic details of the coupling terms, therefore DMMFT is a method generically applicable to fermions, bosons as well as spins, irrespective of the presence of frustrations.
In the following subsections, we begin by reviewing the conventional mean-field approximation in Sec. 2.1. Then, we develop the mean-field equations and self-consistency conditions of DMMFT in parallel to those in conventional MFTs in Sec. 2.2. Following this, in Sec. 2.3, we conduct a comprehensive comparison between the DMMFT and the conventional MFTs. Within this comparison, we identify a hyperparameter, $n_{c}$, that interpolates the DMMFT and the conventional MFTs. Particularly, when $n_{c}=1$, attaining its minimal value, DMMFT becomes equivalent to the conventional MFTs.
@@ 2.1 Conventional Mean-Field Approximation
then, for each extended cluster, $|\phi_{\mathcal{C}^{\prime}_{c}}\rangle=\otimes_{c^{\prime}\in\mathcal{C}^{\prime}_{c}}|\phi_{c^{\prime}}\rangle$. Therefore, Eq.(4), Eq.(5), and Eq.(6) form a closed set of coupled mean-field equations for the conventional MFTs.
The mean-field equations can often be further simplified when systems have additional symmetries. If the partition $\mathcal{C}=\{c\}$ respects some symmetries of the system, there exist symmetry transformations relating the clusters $T_{c^{\prime},c}:c\mapsto c^{\prime}$. A symmetric mean-field Ansatz state can be constructed simply as $|\Phi_{\text{MF}}\rangle=\otimes_{c}T_{c,c_{0}}|\phi_{c_{0}}\rangle$. Particularly, for a translationally invariant state, all $|\phi_{c}\rangle$ can be chosen equal to the same state $|\phi_{c_{0}}\rangle$ in the local Hilbert space $\mathscr{H}_{c_{0}}$, then the coupled mean-field equations for the clusters reduce to a single set of self-consistent mean-field equations for $|\phi_{c_{0}}\rangle$.
@@ 2.2 Density-Matrix Mean-Field Approximation
The entanglement is controlled by $\mathscr{H}_{c}$ than $\mathscr{H}_{\mathcal{C}\backslash c}$ provided $D_{\mathscr{H}_{c}}<D_{\mathscr{H}_{\mathcal{C}\backslash c}}$, where $D_{\mathscr{H}}$ denotes the dimension of the Hilbert space $\mathscr{H}$. According to the purification theorem [59], there exists a state $|\tilde{\Psi}\rangle$ in ~$\tilde{\mathscr{H}}=\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{\tilde{c}}$ such that $\text{Tr}_{\tilde{c}}|\tilde{\Psi}\rangle\langle\tilde{\Psi}|=\tilde{\rho}_{c}$ is equivalent to $\rho_{c}$, and the dimension of the Hilbert subspace ~$\tilde{\mathscr{H}}_{\tilde{c}}$ is bounded by $\tilde{D}^{E}_{c}=\lceil\exp({SE}_{c})\rceil$, where
@@
Instead of Eq.(4), DMMFT assumes that for systems without long-range entanglements, $\tilde{\mathscr{H}}$ can be found as a Hilbert subspace of $\mathscr{H}_{\bar{c}}$ over the extended cluster $\bar{c}$. Mathematically, there exists a homomorphism
@@
Parallel to the mean-field equations for the conventional MFTs, Eqs.(9) – (14) constitute a closed set of coupled mean-field equations for the DMMFT. In the presence of symmetries, the mean-field equations can be further simplified. Specifically, for a translationally invariant target state $|\Phi\rangle$, the reduced DMs of the local clusters are all identical to the same reduced DM $\rho_{c_{0}}$ (so are the projectors). Consequently, the coupled mean-field equations for the clusters reduce to a single set of self-consistent mean-field equations for $\rho_{c_{0}}$. The procedures for implementing the DMMFT algorithm are outlined in the Appendices.
@@ 2.3 Comparison of the Mean-Field Approximations
Comparing conventional MFTs with DMMFT, we find the cut-off parameter $n_{c}$ in Eq.(13) can be regarded as a hyperparameter interpolating the DMMFT and the conventional MFT. To demonstrate this point, we first observe that when $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}=1$, DMMFT is equivalent to a conventional MFT. If the ground state is a product state [Eq.(6)], the local reduced DM of cluster $c$ is $\rho_{c}=|\phi_{c}\rangle\langle\phi_{c}|$. The expectation values of the local observables agree
@@
It is easy to observe that $\rho_{c}$ generically has non-vanishing entanglement entropy whenever $D_{\mathscr{H}_{\mathcal{C}^{\prime}_{c}}}>1$. This is exactly the short-range entanglement captured by the DMMFT beyond the conventional MFTs. Not surprisingly, one expects that the DMMFT may detect short-range entangled topological phases with a properly chosen $n_{c}$. In Sec. 3.1, we demonstrate with the AKLT model that DMMFT can indeed detect the topological ground states in terms of the entanglement spectra without being biased to the symmetry-breaking states.
@@ 3 Applications
In this section, we apply the DMMFT to two demonstrative examples, the AKLT model and the AFHTL, that have been extensively studied in the literatures. The AKLT model was historically proposed as an exactly solvable model to confirm Haldane’s conjecture regarding antiferromagnetic spin-$1$ chains [41, 42, 43]. We choose the AKLT model as our illustrative example not solely because of its exact solvability but primarily due to its intriguing property that the expectations of local moments, and consequently the usual Néel order parameter (alternating magnetization), vanish in the ground states [42, 43], This property challenges the ability of conventional MFTs to detect the orders borne in the ground states of the AKLT model. Moreover, despite proposals suggesting that topological orders in one dimension can be characterized by nonlocal order parameters, such as string order [60], it is still desirable to identify local features that can distinguish topologically ordered states from disordered or trivially ordered states, especially at the mean-field level. In Sec. 3.1, we demonstrate that DMMFT can correctly detect the topological ground states based on the reduced DM and the entanglement spectra. All local observables that can serve as an order parameter can be simply calculated with respect to the reduced DM according to Eq.(15). Nevertheless, the entanglement spectra reveal additional information beyond conventional MFTs, reflecting the quantum nature of orders. It should be emphasized that the topological order in the ground states of the AKLT model is a generic feature of this exotic phase, which is not restricted to the integrable point of the AKLT model, but persists upon perturbations that do not close the gap. The exact solvability of the AKLT model here only serves as a convenient benchmark allowing quantitative comparisons.
The AFHTL serves as a prototypical model representing frustrated magnetisms. Inspired by the pioneering study of Fazekas and Anderson’s [5], it has been proposed that various intriguing phases are potentially hosted by the AFHTL due to the frustrations [6, 61, 62, 63, 64, 14, 65, 66, 67]. In Sec. 3.2, we apply the DMMFT to the simplest case of the AFHTL, which includes only the nearest-neighbor couplings and the easy-axis anisotropy. Even in this simplest case, the ground state of the system exhibits a noncollinear three-sublattice order, reminiscent of the semiclassical Néel state. However, the order parameters are subjected to significant renormalization due to cooperative interactions between quantum fluctuations and frustration-induced noncollinearity [44, 45, 46]. Furthermore, we also report the magnetization of the AFHTL in a longitudinal magnetic field, which has been extensively studied in literatures, both in classical and quantum contexts [51, 52, 53, 54, 55, 56, 57, 47, 48, 49, 50]. Particularly, it has been noted that the quantum fluctuations play non-trivial roles in stabilizing the ordered states in AFHTL [56, 57, 58]. For quantitative comparison, we compute the magnetization curve using the state-of-the-art technique DMRG. Our DMMFT results exhibit good agreement with DMRG. It should be emphasized that the simple AFHTL with only nearest-neighbor couplings and easy-axis anisotropy in a longitudinal external magnetic field makes it a robust candidate for benchmarking purposes because it has been extensively studied using various methods and confirmed in the literatures. Nevertheless, DMMFT is not solely restricted to the simple AFHTL. It can indeed also detect the stripe phases in $J_{1}$-$J_{2}$ AFHTL [14], and the resonating valence bond phases with dimerized spins [5]. We would highlight DMMFT as a generic unbiased approach to characterizing diverse phases with classical as well as quantum orders.